A used car salesman purchases a car from its previous owner at a price of $4,500, although the owner was willing to sell it for as little as $4,000. The salesman later sells the car to a new buyer for $6,000, although that buyer was actually willing to pay up to $6,500.

What percentage of the value created by the trade is captured by the buyer who purchases the car from the used car salesman?


1
Expert's answer
2021-12-21T17:15:15-0500

Total Value created:

"=(4500-4000)+(6000-4500)"

"=2000"

Percentage of value created by the trade that is captured by the buyer,

"=\frac{1500}{2000}\times 100=75"%